### v1

# WildGraph: Realistic Graph-based Trajectory Generation for Wildlife

Abstract.
Trajectory generation is an important task in movement studies; it circumvents the privacy, ethical, and technical challenges of collecting real trajectories from the target population. In particular, real trajectories in the wildlife domain are scarce as a result of ethical and environmental constraints of the collection process. In this paper, we consider the problem of generating long-horizon trajectories, akin to wildlife migration, based on a small set of real samples. We propose a hierarchical approach to learn the global movement characteristics of the real dataset, and recursively refine localized regions. Our solution, WildGraph, discretizes the geographic path into a prototype network of H3111https://www.uber.com/blog/h3/ regions and leverages a recurrent variational auto-encoder to probabilistically generate paths over the regions, based on occupancy. WildGraph successfully generates realistic months long trajectories using a sample size as small as 60. Experiments performed on two wildlife migration datasets demonstrate that our proposed method improves the generalization of the generated trajectories in comparison to existing work, while achieving superior or comparable performance in several benchmark metrics. Our code is published on the following repository: https://github.com/aliwister/wildgraph.
small data, data mining, trajectory generation, wildlife movement
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Trajectory data provide invaluable insights into the patterns and behaviors of movement, enabling a deeper understanding of the underlying mechanisms driving movement across various domains, including human, vehicle, and wildlife. In particular, wildlife trajectories are utilized by scientists across different fields to promote wildlife conservation (Wall et al., 2014), anti-poaching efforts (Park et al., 2015), climate change studies (McRae et al., 2008), human-wildlife conflict (Buchholtz et al., 2020), and to help guide policy towards wildlife preservation (Tomkiewicz et al., 2010). For example, trajectories help identify where animal movement may intersect with known poaching activity, and hence promote anti-poaching efforts. Similarly shifting in trajectory
However, the collection process of real trajectory data may be associated with various challenges. Wildlife researchers have cited ethical and technical constraints involved in wildlife trajectory data collection (Dore et al., 2020). In particular, data is traditionally collected using devices attached directly to the subject, which requires capturing and possibly sedating it. The devices are further constrained by the size of the animal, and may be prone to intermittent or permanent failures, thus limiting the availability of trajectory data.
(a) WildGraph
(b) Real
(c) GAN
Figure 1. WildGraph vs. Real vs. GAN-generated Trajectories
In order to overcome challenges associated with collecting real movement data, trajectory generation has been considered in various domains, including human mobility (Long et al., 2023; Yuan et al., 2022; Rossi et al., 2021b), vehicle mobility (Cao and Li, 2021; Rossi et al., 2021a), as well as wildlife (Al-Lawati et al., 2023; Technitis et al., 2015) movement studies. Ideally, we want the generated set of trajectories to have the same characteristics (measured by, e.g. Hausdorff Distance or correlation coefficient) of a set of real trajectories, and generates original samples that preserve the patterns of the real trajectories (e.g., clusters or distance similarity). For example, in Figure 1, it can be visually verified that WildGraph generates much better trajectories, than using a Generative Adversarial Network (GAN) for this problem. The generated trajectories provide valuable insight in experimental studies, scenario planning, and risk assessment, where researchers can simulate the effects of polices and environmental decisions, and thereby improve their robustness and generalization (Beach and Clark, 2015; Catano et al., 2015).
However, unlike human and vehicle trajectories, wildlife trajectory generation presents a unique set of challenges. First is the small data problem: real trajectory datasets are sparse and include a small number of subjects. Second, the length of each trajectory tend to span several months, or years (e.g. in that long timespan, they may vary due to confounding factors, weather patterns, extreme events that are outliers, etc.) to provide practical value. Finally, often there are no established network information, e.g. roads, which wildlife may follow, (although in some land animals there may be tracks on the ground trailed by the migrations, these are not rigid and animals often stray from them to create new trails/paths easily) thus increasing the complexity of the generation task.
Despite these challenges, interestingly, simple deep learning based techniques, such as Variational Auto Encoders (VAEs) have demonstrated promising results in generating long-horizon trajectories using a small set of real training samples (Chen et al., 2021; Al-Lawati et al., 2023). VAEs accomplish this by mapping trajectories in a probabilistic manner to a latent space, and teaching the decoder to reconstruct the encoded trajectory from the latent representation. Hence, new trajectories that have similar properties as the real trajectory may be generated by sampling the latent space.
However, while samples generated from VAEs appear plausible, they fail to generalize to paths that have not been observed during training. In addition, point selection along the trajectory by VAEs may not be geographically valid (see Fig 6(e)). For example, just as vehicles may follow a specific road network, and may stop at rest areas and gas stations, wildlife tend to congregate around water bodies, and vegetation, but not in the middle of the ocean.
In this paper, we address these challenges by constructing a prototype network of regions for wildlife movement that models trajectories as a path across the network. We utilize the set of known trajectories to construct the global regions, and recursively refined localized regions, aka nodes, into multiple granular nodes. This helps preserve the global direction of movement, and at the same time captures the local properties of the generated trajectories. By adjusting the region sizes of the network to limit the selection of potential points from a particular region, we improve the generalization of the generated trajectories, and provide a better outcome across multiple benchmarks.
Our proposed framework, WildGraph, models trajectories, real or generated, as paths across this network of regions. WildGraph leverages the H3 geographic library, to construct the network. H3 is an indexing system for map regions that has seen increasing use across trajectory-related research (Han et al., 2021; Musleh and Mokbel, 2023). It provides a standard, unbiased representation of each region, i.e. the choice of the regions and their sizes are not affected by any given trajectory and does not overfit the training data or a small subset of it.
Once each known trajectory is encoded as a path over the network, we process it using a Variational Recurrent Network (VRN), which trains a model that generates a dictionary of latent spaces for each region, as depicted in Figure 3. The details of our framework is discussed in further details in Section 3.
The contributions of this work can be summarized as follows:
  • We propose WildGraph: a framework that generates realistic long-horizon wildlife trajectories based on a sparse set of real trajectories.
  • We contribute a novel approach in prototyping trajectories as paths across a network, and demonstrate its effectiveness in improving inference.
  • We conduct multiple experiments against two wildlife datasets, and benchmark our methods against our earlier work, WildGEN (Al-Lawati et al., 2023), a Variational Auto-encoder (VAE) based approach, Generative Adversarial Network (GAN) based approach, and a Transformers-based approach based on the prototype network.
To the best of our knowledge, this paper along with our previous (unpublished) workshop paper: WildGEN (Al-Lawati et al., 2023) are the first to propose deep-learning based trajectory generation for wildlife movement. Specifically, WildGEN, demonstrated that deep learning approaches yield significantly more realistic trajectories than traditional methods, such as Correlated Random Walk (CRW) (Renshaw and Henderson, 1981) and Levy Walk (Wosniack et al., 2017). This study builds upon our previous work by refining the point selection process along trajectories, as assessed by the similarity between generated trajectories and real trajectories in our test set.
(a) *
Coarser zoom level
Finer zoom level
(b) *
(a) *
[.23]Refer to caption [.23]Refer to caption
Figure 2. H3 Regions of different zoom levels covering the set of known points
2. Related Work
In this section, we discuss some recent works that are related to the methods and solutions we present in our paper, from different research areas.
Trajectory Generation. In the pre-deep-learning era, trajectory generation was relatively subject specific. In ecology, researchers used methods such as Levy Walk (Wosniack et al., 2017), Correlated Random Walk (CRW) (Technitis et al., 2015), and other velocity based methods (Technitis et al., 2015; Ahearn et al., 2017), that rely predominantly on the subject (animal) considered. The advent of deep-learning-based generative models such as VAEs and Generative Adversarial Networks (GANs) resulted in some of the earliest attempts at using these approaches for trajectory generation. Chen, et al. (Chen et al., 2021) benchmarked GANs and VAEs for the task of generation, and found that VAEs provide more consistent results. Our earlier work: WildGEN, was among the first methods to address the problem of trajectory generation for wildlife (Al-Lawati et al., 2023). We utilized VAEs with added layers of post-processing that mitigated some of the challenges discussed above. However, much like other VAE-based methods, it suffers from limited generalization capabilities.
Most recently, Neural Temporal Processes were proposed for the task of generation in human mobility (Wan and Dodge, 2023). This approach overcomes challenges in dealing with irregularly structured data, which may have various applications in wildlife. However, in our case, the datasets comprise of regularly sampled data.
Trajectory Similarity. Trajectory similarity is a complementary problem to trajectory generation since the goal of generation is to produce a set of trajectories similar to a set of known trajectories. An interesting approach to the problem was considered by Yao, et al. (Yao et al., 2022), where a region is recursively subdivided to capture the region based similarity between compared trajectories. Similar to our approach, they recursively split regions into smaller regions to create a balanced graph. Next, they utilized a graph-based attention layer to model the embedding of each node. In comparison, our approach utilizes node2vec (Grover and Leskovec, 2016) to overcome the over-parameterization in our problem given the disparity in the training sample sizes. Finally, since the trajectory is considered as a whole in the trajectory similarity problem, a snapshot of the graph may be processed based on the Normalized Lagrangian of the adjacency and degree matrices. A time-aware approach to generate similar graph snapshots at each step in a generation problem would render the problem considerably inefficient. Moreover, it will fail in the presence of cycles in the graph because it cannot determine the correct sequence of nodes in the trajectory.
Graph Generation. Early generative graph models were based on adjacency matrix generation using encoder-decoder structures (You et al., 2018; Simonovsky and Komodakis, 2018). However, these approaches fail to capture the dynamic nature of spatio-temporal graphs. Recently, models such as STGD-VAEs (Du et al., 2022), and STGEN (Ling et al., 2023) were proposed to address these limitation. STGD-VAEs factor graphs into dependent ordinal values and jointly optimize the objective. On the other hand, STGEN utilizes the distribution of spatiotemporal walks to adversarially learn the characteristics of the graph. However, multi-objective approaches appear to be a poor fit for this problem, given the small sample sizes, and long horizon. We validate this in our experiments and demonstrate the GANs provide poor results on all metrics (see  1).
Sequence Modeling. Sequence-to-Sequence models have been used for various tasks in wildlife trajectory studies, including prediction (Feng and Gordon, 2019), imputation (Li et al., 2021) and interpolation (Wan and Dodge, 2023). They identify implicit patterns by using LSTM or RNN memory trained with new observations in an auto-regressive manner (Park et al., 2018). However, the multi-variate nature of spatial data often results in a quick deterioration of predictions on tasks involving long horizon prediction (Sagheer and Kotb, 2019), particularly in small data settings. To address this, Li, et al., (Li et al., 2021) attempted to simplify the problem to a prediction of step size instead of location, but this approach offers very limited usefulness. Our work adopts a more practical approach by leveraging the properties of the H3 spatial library to represent geospatial regions, where each region is encoded with a string representing its address encoded in the library, thus reducing the problem of generating a path to one in a single dimension.
In addition, encoding each region’s connection (i.e., the probability of whether a bird will fly from one region to another) with other regions provides additional context to the problem, and enables us to generate long horizon trajectories based on a small set of training samples.
3. Preliminaries
3.1. Problem Statement
Given a set of $n$ real trajectories in an area $\mathcal{A}$, each consisting of $m$ points, the objective is to produce a set of synthetic trajectories that maximize the following:
  1. (1)
    (Path Similarity) similarity between the points of the synthetic trajectories to that of the real trajectories, and,
  2. (2)
    (Likeness) the alignment of the distribution of the synthetic trajectories with that of a (test) set of real trajectories.
Formally, the task is to learn a model to generate synthetic trajectories that maximize the similarity between a set of held-out test-set comprising of $k$ real trajectories, denoted as $T_{1},T_{2},\ldots,T_{k}$, each comprising $m$ spatio-temporal points, where $T_{i}=(p_{1}^{i},p_{2}^{i},\ldots,p_{m}^{i})$, with $p_{j}^{i}$ representing latitude ($y$) and longitude ($x$) values. The objective is to find a corresponding set $S={S_{1},S_{2},\ldots,S_{t}}$ of synthetic trajectories that best matches the (test set of) real trajectories.
To achieve this, we aim to maximize the overall similarity between $T$ and $S$, denoted as:
$\max\text{ Similarity}(T,S)+\max\text{Likeness}(T,S)$
Here, the similarity measure encompasses both Path similarity, and Likeness. Note that using a held-out set to compare with the generated trajectories, post-training, ensures robustness and generalization of the proposed solution. Further details on the evaluation metrics utilized are provided in Section 3.5.
3.2. Overview of WildGraph
Figure 3 demonstrates the overall structure of WildGraph. The training phase proceeds in the following way:
  • WildGraph accepts a sequence of training samples of known trajectories, each of which consist of $m$ points. All of these points are used in the Hierarchical Network Generator, which implements a prototype graph ($\mathcal{G}$) based on them. A node represents a region of $\mathcal{G}$, while the edges represent transitions from one node to another as observed in a given trajectory.
  • A pre-training of the nodes of $\mathcal{G}$ is performed based on the connections identified. We utilize node2vec to generate representations of each node based on the input and output connections. The pre-training allows us to encode the relationships between the nodes and provides a representation learning of the transitions in the network, as well reduces the dimensions of the embeddings, compared to structures such as Bag of words (BOW).
  • The Variational Recurrent Network (VRN) aims to learn the empirical probability distribution of all states, i.e. learning the transition kernel, in order to generate trajectories following this distribution in an auto-regressive manner. Post-training, the output distribution of each node is aggregated in a latent dictionary (described below) that is used later during the generation process.
Once training is completed, the process of generating new trajectories utilizes the latent dictionary to stochastically select the next region based on the current region. The latent dictionary is a multi-valued dictionary that collects latent values ($z$) that correspond to a given state. The generation process is repeated until the desired sequence length (input parameter) is reached. For each region, an occupancy based sampling converts the region into a point. The occupancy sampler divides the entire area $\mathcal{A}$ into smaller regions, i.e. dots, and builds a discrete heatmap on the entire area. The process of converting a region to a point involves sampling a dot (a small region) from the corresponding region based on the heatmap, and randomly sampling a point from the dot.
3.3. H3 Grid System
As mentioned above, we utilize the H3 geospatial indexing system. H3 is an open-source library for creating and managing hierarchical spatial indexes. The core concept of H3 is the use of hexagonal grids that partition the map into smaller spatial units, at different zoom levels. The developers cite various advantages that include: (1) capacity to model large spatial regions and have better descriptive ability of spherical surfaces; (2) provides a single string-based address (e.g. 821fa7fffffffff) that encompasses a geographic region of points; (3) built-in hierarchical description of a region at 14 zoom levels (Uber Technologies, Inc., cess).
Note that H3 divides up the earth’s surface into a set of hexagonal regions that are fixed once the zoom level of the hexagon is fixed; however one can choose smaller-sized hexagons to zoom into a region if so required.
Figure 2 demonstrates points in the same $\mathcal{A}$ covered by regions using different zoom levels of H3. As mentioned above, each hexagon in either image has a unique string-based address. Our framework utilizes several of the features and benefits provided by H3.
4. Methodology
Figure 3. WildGraph Framework
Figure 3 depicts the general architecture of WildGraph. The framework is composed of five main components:
  • Hierarchical prototype Network Generator (HNG)
  • Graph-based Embedding Layer
  • Variational Recurrent Network (VRN)
  • Latent Dictionary
  • Occupancy Sampler
We describe each component in further detail in following subsections.
4.1. Hierarchical Network Generator
(a) *
Regions divided into sub-regions
A graph of transitions is captured and used to train a graph embedder
(b) *
(a) *
[.33]Refer to caption [.66]Refer to caption
Figure 4. Hierarchical Network Generator. The red points are the observed data points of where the animal was whereas the black lines are the linear interpolations connecting them.
The Hierarchical Network Generator (HNG) takes the set of known trajectory points and generates a prototype network graph ($\mathcal{G}$) that represents $\mathcal{A}$. Initially, a low zoom level is used to discretize $\mathcal{A}$, which promotes generalization. However, for each region that encompasses more than a single point, the diameter of the region is calculated and used as basis to discretize it into a higher zoom level as depicted in Figure 4. This processing of localizing regions helps preclude unlikely transitions in a generated trajectory.
The diameter is defined as the maximum distance between two points in a given region. Given a set of points $(p_{1},p_{2},....,p_{n})$ in a given region ($h3$), the diameter is:
$diameter(h3)=\max||p_{i}-p_{j}||\hskip 14.22636pt\forall p_{i},p_{j}\in h3$
where —— · —— denotes the Euclidean norm (or distance). In the event $diameter(h3)>r$, for some split threshold parameter ($r$), the region is split, and all points $(p_{1},p_{2},....,p_{n})\in h3$ are reassigned more granular H3 regions. This process is repeated recursively as captured in Algorithm 1.
Note that excessive zooming results in a large network that consists of too many nodes, which may limit the generalization of the model, and increase its complexity. The optimal threshold $r$ is learned from the training data. It depends on properties of the dataset, and may be influenced by the resulting average density of the graph. In the experimental section, we empirically investigate how the choice of $r$ affects the evaluation metrics.
Next $\mathcal{G}$ is constructed for observed transitions between regions as depicted in Figure 4
1:  Initialize $zoom\leftarrow 2$ //Starting H3 regions at this zoom level
2:  Initialize $P\leftarrow(p_{1},\ldots,p_{i\times m})$
3:  $H\leftarrow Assign\_H3(P)$ //Assign H3 address to each point
4:  while $\forall h3\in H3\mid diameter(h3)>r$ do
5:     $zoom\leftarrow zoom+1$
6:     $P_{s}\leftarrow P[h3]$ //Filter points that need granular H3 regions
7:     $H_{s}\leftarrow Assign\_H3(P_{s})$
8:     Update $P$ with $P_{s}$ //Propagate updates
9:     Update $H$ with $H_{s}$ //Propagate updates
10:  end while
11:  return  $P,H$
Algorithm 1 Regionalize Trajectories([$T_{1},\dots,T_{n}$], m)
The final step involves creating the edge list of the network. Algorithm 1 regionalizes each trajectory from a sequence of points to a sequence of regions:
$T_{1}=(h3_{1}^{1},h3_{2}^{2},\ldots)$
$T_{2}=(h3_{2}^{1},h3_{2}^{2},\ldots)$
[…]
The edges of $\mathcal{G}$ are the set of all $(h3_{i}^{k},h3_{j}^{k})$ pairs in each trajectory. While $\mathcal{G}$ is a directed graph, no weights are assigned to the edges.
4.2. Network Graph Embedding
A pretraining of $\mathcal{G}$ is performed to learn its characteristics and create embeddings for each node. For simplicity, we use a standard graph-based skip-gram embedding algorithm, namely node2vec to generate the embedding for each state. Node2vec is network representation learning technique to generate low-dimensional vector representations, i.e. embeddings, for the nodes in a graph. It performs a random walk, enabling it to efficiently capture both local and global structural information within a graph (Grover and Leskovec, 2016).
Node2vec employs a two-step random walk procedure to explore the graph space. First, it performs short random walks to capture local neighborhood information around each node. These walks balance between breadth-first and depth-first exploration, controlled by two parameters, $p$ and $q$, which dictate the likelihood of revisiting recent nodes or exploring new ones.
Given, the long-horizon nature of our trajectories, we select a higher value of $q$ to explore more distant nodes. The resulting embeddings learned by node2Vec encode rich structural information about nodes. Node embeddings also significantly reduce the dimensionality of the problem and facilitate out of policy exploration: transitions that are not observed in the real dataset, but are potentially realistic transitions.
4.3. Variational Recurrent Network
The Variational Recurrent Network (VRN) uses the embedded sequences of trajectories to train the model by processing them sequentially. It combines elements of recurrent neural networks (RNNs) with variational inference techniques to model the sequential decision-making under uncertainty. Analogous to conventional RNNs, the VRN incorporates recurrent connections that allow it to maintain a hidden state representing the network’s memory of past observations and actions. This hidden state evolves over time as the network processes the input sequentially.
The VRN is trained using a combination of unsupervised learning and recurrence techniques. Unsupervised learning is used to train the network to predict the next observation or action given the current observation and hidden state. The utilized observation is the current belief vector over the states, obtained as the raw output of the VRN. Given the recurrence nature of VRN, the belief vector is trained on the history of the previous states. This approach may be viewed as learning a transition kernel from the current state to future states while considering the history of the trajectory. Post-training, the hidden state, the output of the bottleneck layer in the autoencoder, is statistically treated to reflect the possible transitions per the training trajectories, though it is represented in a compressed space.
The VRN utilizes an encoder-decoder model. Given a pair of region-mapped and graph embedded points $(h3_{i},h3_{i+1})$, the encoder maps $h3_{i+1}$ to the latent space representation $z_{i+1}$. The decoder attempts to reconstruct the input vector $h3_{i+1}$ from a concatenation of $z_{i+1}$ with $h3_{i}$, modified with positional encoding. The converged $z_{i+1}$ after training is captured and added to the multi-dictionary of $h3_{i}$.
During inference, the VRN makes sequential decisions by iteratively updating its hidden state based on observed inputs and selecting actions according to its learned policy. The network’s belief state is updated, allowing it to maintain a probabilistic estimate of the environment’s true state given the past states.
4.3.1. Positional Encoding
Positional encoding (PE) assigns unique representations to each position in the sequence. This allows the model to differentiate between regions based on their position in the sequence. For example, a bird may remain in a given region for several days before proceeding to the next region. The PE enables the model to learn from the sequential context effectively. We use a simple approach by concatenating the raw index of the position to the input of the decoder. This helps the network to associate certain transitions with the position, while avoiding the risk that the VRN may fail to learn a more sophisticated method given the small data nature of the problem.
4.4. Latent Dictionary
After training, the representation of each possible state $h3$ is passed through the encoder to find the corresponding latent space realization $z$. Using the frequentist approach, for each node $h3$, the possible transitions are collected in a dictionary and then replaced by their corresponding latent space realizations. When a certain node appears as the current state in our VRN, we sample uniformly from the latent dictionary of this node. This is akin to sampling the stationary next state based on the possible transitions in the training set, but in latent space. However, these stationary transitions are combined with the belief vector, thus learning the non-stationary transitions, while taking into consideration all previously generated states.
Formally, the dictionary $D$ can be represented as a mapping for all nodes $(k_{1},\ldots,k_{s})$ in $\mathcal{G}$:
$\{k_{1},\ldots,k_{s}\}:\{(l_{1}^{j_{1}}),\ldots,(l_{s}^{j_{s}})\}$
For all pairs of transitions between nodes in the training set: $(k_{i},k_{j})$, the dictionary is updated: $D^{\prime}=D\cup\{k_{i}:(l_{j})\}$ where $l_{j}=enc(k_{j})$, and $enc$ the trained encoder from the VRN.
4.5. Occupancy Sampler
The occupancy sampler chooses a dot (a very small region) and then from the dot, it chooses a point at random. The sampling is conducted based on the heatmap of the region as depicted in Figure 5. The sampler proceeds as follows:
  1. (1)
    Split the region into dots, which in our case are H3 regions with a high zoom that fall within the region.
  2. (2)
    Based on the distribution of real points inside the regions, sample a dot.
  3. (3)
    Sample a point that lies inside the dot randomly.
The intuition behind using a two-step hierarchical sampler (i.e., going to dots and then choosing points) as opposed to choosing points from a heatmap directly is as follows. If we represented a region as a set of points and used a heatmap to choose the points directly, then we would be choosing points from existing trajectories with high probability. By introducing a small region that we select based on the heatmap and randomizing the choice of the point from the dot, we introduce a level of generalization that is useful in introducing additional stochasticity in point selection. An added advantage of this approach is it helps ensure the points selected have geographic validity, i.e., are close enough to actual points because of how we choose the output points based on a dot, which is a very small region. This can been visually confirmed from the generated paths in Figure 6(a), where WildGraph trajectories, much like real trajectories, almost always avoiding congregating in the middle of the sea, while this is not the case with some of the benchmark methods, even competitive methods such as WildGEN.
Figure 5. Occupancy Sampling of dots based on heatmap
5. Experiments
5.1. Baselines
In order to assess the performance of our framework, we compare the results obtained to those from existing methods for animal trajectory generation. As discussed earlier, CRW-based and other traditional motion methods fail at generating long-horizon trajectories for tasks such as migration. This was demonstrated in our workshop paper (Al-Lawati et al., 2023).
As such we restrict our evaluation to deep-learning based baselines that appear to be plausible for the task of long-horizon trajectory generation, namely:
  • Variational Auto-encoder (VAE)
  • WildGEN
  • Generative Adversarial Network (GAN)
  • Transformers
In the following subsections, we briefly introduce each method.
5.1.1. VAEs
As discussed earlier, VAEs are a supervised training model that map inputs into a latent space while minimizing the distributional divergence between the input and the output. The latent space is subsequently sampled to generate original samples.
5.1.2. WildGEN
WildGEN is a model based on VAEs that performs post-processing on the generated trajectories. Two forms of post-processing are applied:
  • Smoothing: utilizes a Savitzky-Golay smoothing filter (Press and Teukolsky, 1990) to reduce excessive wandering.
  • Minimum Bounding Region: drops any trajectories that venture out of a minimum bounding polygon of the set of known points.
5.1.3. GANs
Generative Adversarial Network (GAN) is a model that consist of two neural networks, the generator and the discriminator, which engage in a minimax game. The generator aims to produce realistic synthetic data samples from random noise, while the discriminator learns to distinguish between real and generated samples. Through adversarial training, GANs progressively improve both the generator’s ability to create realistic data and the discriminator’s proficiency in discerning real from fake. This adversarial process drives the generator to generate increasingly realistic samples (Goodfellow et al., 2020).
5.1.4. Transformer
Transformers are networks that have the capacity to auto-regressively generate sequences of tokens based on attending to previous tokens. The HNG facilitates the use of H3 word-based language transformers, such that a trajectory is a sentence that consists of $m$ words: each of which is an H3 address.
5.2. Evaluation Metrics
Our selection of evaluation metrics is based on the objectives of trajectory generation, which we reiterate here for convenience:
  1. (1)
    (Path Similarity) How similar are the points of the synthetic trajectories to the real trajectories?
  2. (2)
    (Likeness) How well do the synthetic trajectories align with the real trajectories in terms of areas of concentration?
In order to measure Path Similarity, we use three distance metrics: Hausdorff Distance, which measures the steepest gap, and Dynamic Time Warping (DTW) which measures the spatial similarity over the whole path. In addition, we also measure the Final Displacement Error (FDE) which we define as the gap between the final positions of a generated trajectory and the nearest real trajectory: $\text{FDE}=\|p_{t}^{m}-\hat{p}_{s}^{m}\|$.
On the other hand, we measure Likeness using Pearson Correlation Coefficient (r-Coeff) and Chi-squared ($chi^{2}$) similarity, between the generated trajectories and the held-out real trajectory set.
For each generated trajectory, we compare it to every trajectory from the test set, and use the closest match. The reported distance metric for the experiment is the mean of these values over a fixed number of generated trajectories.
5.2.1. Coverage
In addition, we introduce a new evaluation metric called coverage (cov) to measure how representative the generated trajectories are of the test set of real trajectories. This measure, which is captured for each distance evaluation metric above, quantifies the percentage of the held-out real data that provided the measurement from the generated trajectories. For example, given generated trajectory $S_{i}$, if the best Hausdorff or DTW measure was observed when compared to a real trajectory $T_{j}$, we add $T_{j}$ in the coverage list. This measure quantifies the fraction of trajectories added in the coverage list of each distance metric.
5.3. Datasets
We utilize two dataset sets of migratory birds obtained from Movebank [^2] https://www.movebank.org. Both datasets were processed using Moveapps (Kölzsch et al., 2022) to limit observations to one per day for each subject. For both datasets, any subject that had significant missing readings were dropped. Subjects that were tracked over multiple years were split into multiple samples. Simple data interpolation was utilized to fill in any remaining isolated gaps in the trajectories. The two datasets are:
  • The Western Palearctic greater white-fronted geese dataset (geese dataset) (Kölzsch et al., 2019), which was clipped to a window between March 1 to September 1 (185 days) to capture northward migration. The resultant dataset consists of 60 trajectories over a period of 185 days.
  • The White Stork Ciconia ciconia dataset (stork dataset) (Carlson et al., 2021), which was clipped to a window between August 15 to November 1 (79 days) to capture southward migration. The resultant dataset consists of 77 trajectories over a period of 79 days.
5.4. Experimental Settings
Preliminary tests were performed to evaluate the convergence properties of each method on a fixed learning rate of $31e-$ 3 using Adam optimizer. For the GAN experiment, additional decay rate parameters were used to induce convergence.
Each experiment used 5-fold cross validation to generalize given the small data nature of the problem. All reported values are the average of 5 identical experiments. This choice was made to normalize the randomness of the generation process, and report more consistent results.
The experiments were conducted utilizing a single 80GB A100 Nvidia GPU paired with 8 CPUs, operating on Python 3.8, PyTorch 2.1.1, and CUDA 11.8.
5.4.1. Distributional Settings
In order to measure the likeness (r-Coeff and $chi^{2}$), we utilize the discrete cluster distributions between the generated points and the test set. In order to create the clusters, the optimal cluster size for K-means was pre-calculated and used as input in the distributional likeness measurements. The cluster sizes were determined using a Silhouette graph and selected visually. Specifically, the cluster sizes for the geese and stork datasets are 16 and 14, respectively.
5.5. Results
Table 1. Test Results for the geese and stork datasets using WildGraph and other benchmarks using different measurement
Dataset Method Hausdorff DTW FDE r-Coefficient $chi^{2}$
val cov val cov val cov
Geese VAE 7.45 0.57 57.67 0.68 3.81 0.80 0.47 1588
GAN 25.25 0.30 196.8 0.27 11.87 0.44 -0.47 11894
WildGEN 7.12 0.72 48.60 0.79 3.2 0.86 0.79 772
WildGraph 9.38 0.93 67.87 0.88 2.61 0.78 0.79 725
WildGraph-Transformer 22.01 0.90 225.73 0.85 14.26 0.82 0.53 2232
Stork VAE 7.74 0.64 41.1 0.58 3.86 0.53 0.56 676
GAN 14.54 0.19 58.97 0.17 3.96 0.34 0.61 1501
WildGEN 5.55 0.61 28.41 0.58 3.12 0.61 0.76 611
WildGraph 7.29 0.80 40.16 0.77 3.74 0.69 0.8 357
WildGraph-Transformer 16.07 0.64 92.92 0.62 9.45 0.51 0.62 631
Table 1 lists the results obtained for WildGraph and the other benchmarks on the metrics described above. WildGraph achieves superior overall performance on most generalization measurements, including coverage results, r-Coeff, and $chi^{2}$. On the distance metrics, WildGraph reports comparable results to VAE and WildGEN on both datasets.
5.5.1. Split Threshold
As discussed in section 4.1, here, we analyze the affect of the split threshold on generalization. Figure 6 depicts how coverage values, and r-Coeff is affected for different values of the split threshold.
Figure 6. Generalization as a function of split threshold
Based on the results of Figure 6, a split threshold (r) value of 0.25 is used in all of our experiments, as it provides the highest r-Coeff.
5.6. Ablation Study
In the ablation study, we want to measure the effects of the HNG and the node2vec embeddings on the measured metrics. This helps explain how each component contributes to the overall structure.
Table 2. Ablation Study on the Geese Dataset
Method Haus. DTW FDE r-Coef
val cov val cov val cov
Uniform H3 / Coarse 9.65 0.85 74.1 0.80 3.90 0.88 0.63
Uniform H3 / Granular 10.58 0.83 84.49 0.87 3.47 0.67 0.75
BOW (no node2vec) 37.5 0.08 224.16 0.08 0.74 0.17 0.73
No PE 9.06 0.88 69.46 0.83 3.6 0.80 0.72
WildGraph 9.38 0.80 67.87 0.88 2.61 0.78 0.79
(a) WildGraph
(b) Real Samples
(c) WildGEN
(d) WildGraph-Transformers
(e) VAE
(f) GAN
Figure 7. Random Real and Generated Trajectories from WildGraph and the benchmarks
As evident from Table 2, node embeddings emerge as the most impactful component within the WildGraph framework based on the observed results. Using BOW instead provided poor results possibly due to both its high dimensionality and the absence of node relationships in the embeddings.
In addition, by running experiments with fixed region sizes at both high zoom (Granular) and low zoom (Coarse) levels, it can be observed that the HNG is a key component that improves trajectory similarity specifically, but also the likeness on most evaluation metrics.
5.7. Qualitative Analysis
Figure 7 depicts the performance of WildGraph in comparison to other benchmarks. A preliminary visual examination reveals that WildGraph trajectories exhibit more realistic paths. Unlike WildGEN and VAEs, and similar to the real samples, trajectories do not cluster in the middle of the sea. It is also observed that GAN and Transformer-based approaches provided poor results, which is expected given the small data nature of the problem.
6. Conclusions
In this paper, we study the problem of trajectory generation over a long horizon, with a focus on applications in wildlife in a small data setting. Existing methods fail to consider the underlying network that describes that movement. In order to solve this problem, we introduce a hierarchical prototype network graph that captures the global movement characteristics and recursively localizes regions. We propose a novel Variational Recurrent Network to capture the transitions and feed them into a latent dictionary. The effective use of the H3 library, and node2vec embeddings has contributed to the success of the WildGraph framework, as measured in the experimental results and ablation study.
This study adds to the body of literature where graph based solutions can effectively address unique problems. Ours is the first one to show how one can generate trajectories for wildlife movement. In our case, the small data nature of wildlife datasets combined with wide horizon predictions presented significant challenges. Our results demonstrate the effectiveness of our proposed approaches that utilize specially designed encoder-decoder networks. In the near future, we will explore how our generated trajectories can be used for data augmentation and training larger scale models in small data settings.